A 6000 kg rocket is set for firing.If the exhaust speed is 1000 m/s, how much gas must be ejected each second to supply the thrust needed to overcome the weight of the rocket???

Grade:11

1 Answers

Maithri
6 years ago
m=6000kgu=1000m/sF(thrust)=u(dm/dt)Weight=mg=6000×10=60000f(thrust) = weightU(dm/dt)=mg1000(dm/dt)=60000(dm/dt)=60kg/s required to overcome the weight of the rocket.